### v2

@@ 5.1.2 Prompt Analysis
The results in Table 4 highlight the positive impact of combining Automatic-Chain-of-Though (ACoT) reasoning with Few-Shot (FS) examples on the performance of LLAMA 3.1 8B across different modalities in the Assembly101-O dataset. In all prompting scenarios, the system prompt is written as: {systemframe} System: I am going to provide an input sequence that represents a sequence of actions. Your task is to predict the next action of the last sequence based on the patterns observed in the provided input. Limit yourself to only answer with the predicted sequence, and follow the same format given as input.
@@
In Zero-Shot (ZS) settings, the instruction prompt is presented as: {systemframe} System: Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
@@
[…]
@@
When comparing ZS, FS, and ACoT prompting methods, ACoT outperforms the rest of the prompting techniques (Table 5), particularly when using the pre-trained model (Base), across both textual and numerical tasks. In the ACoT approach, automatic ACoT reasoning is employed, where the LLM is queried twice: first, to generate an internal reasoning step for the answer, and then to use this reasoning, combined with a few examples (as in FS), to derive the final answer. The first query is:
@@
ACoT, in contrast, excels in these scenarios because it prompts the model to break down the task into a series of logical intermediate steps before concluding. This approach mirrors the “self-explanation” cognitive strategy, where breaking down and explaining each step improves problem-solving skills. This explains why ACoT significantly outperforms ZS and FS, especially in tasks that demand nuanced understanding and intermediate reasoning.
@@ 5.1.4 Speed Analysis
Table 6 presents a speed test comparison of different prompting methods—ZS, FS, ACoT—using LLAMA 3.1 on the Assembly101-O dataset. The results indicate that ZS is the fastest method, with an average processing speed of 0.208 seconds per sample. FS is slightly slower at 0.216 seconds per sample, representing a marginal 3.8% rise in processing time compared to ZS. ACoT, while offering superior accuracy and reasoning ability, is the slowest, taking 0.315 seconds per sample, which is a 51.4% increase in speed compared to ZS. These differences in speed can be attributed to the inherent complexity of each method. ZS and FS prompt the model to generate answers directly, with FS introducing a minimal overhead due to the additional context provided by the few examples. On the other hand, ACoT requires the model to engage in more complex, step-by-step reasoning, which naturally demands more computational resources and time. While ACoT’s performance benefits are clear, these come at the cost of slower processing, which could be a consideration in time-sensitive applications. Thus, the choice of prompting method should balance the trade-off between speed and output quality based on the task’s specific requirements.
@@ 5.2 Aggregation Strategy
We evaluated three distinct frame aggregation strategies (Fig. 6) aimed at refining per-frame predictions into consistent action sequences, effectively reducing noise and enhancing stability. This results in a cleaner input for the anticipator module.
@@
Each aggregation strategy’s performance was evaluated using the Levenshtein Similarity between the predicted and ground truth action sequence, which derives from the Levenshtein distance.
The similarity between two action sequences (i.e procedures) A and B is defined as:
Levenshtein $\text{Levenshtein Similarity}(A,B)=1-\frac{d_{Lev}(A,B)}{\max(len(A),len(B))}$ (6)
@@
Non-Overlapping Mode Aggregation (NOMA)
The first aggregation strategy we propose (Figure 4) consists of a two-step process. In the first step, we calculate the mode for each non-overlapping sliding window of frames of a given length $L$. Because the windows are non-overlapping, this approach introduces a slight delay in processing, which is a trade-off for improved prediction stability. The mode, representing the most frequent action label within that window, replaces all predictions in that window, ensuring consistency across the frames and removing the noise of smaller portions wrongly predicted. In the second step, we apply the elimination of successive duplicates to remove consecutive repeated action labels, which prevents redundant predictions. This cleaned output serves as the input to the anticipation module, ultimately aiming to produce a more stable and accurate sequence of action predictions.
Overlapping Mode Aggregation (OMA)
The second aggregation strategy we propose (Fig. 4) employs a sliding window approach with a stride of $1$ frame, allowing for overlapping portions of frames. In this method, we evaluate the mode within the overlapping window that includes the last frame and the previous $L-1$ frames. The prediction of the last frame is then substituted with this mode, which captures the most frequent action label in that localized area. This method is designed with the intent to reduce delay, as it processes frames more dynamically compared to non-overlapping windows.
Overlapping Centered Mode Aggregation (OCMA)
@@ 6 Conclusion
Real-time processing constraints can present challenges, particularly when incorporating Large Language Models (LLMs), which may introduce latency during inference. Our focus is maintaining accuracy in an online setting, rather than optimizing for real-time performance. Nonetheless, real-time optimization remains a promising direction for future work.
While symbolic reasoning simplifies certain aspects of action prediction, it may not fully capture the complex semantic relationships between actions. Future research could explore hybrid approaches that integrate symbolic reasoning with richer representations.
Lastly, our evaluation metrics primarily emphasize action recognition and anticipation accuracy. However, other aspects, such as robustness to noise, the ability to handle ambiguous scenarios, and the interpretability of predictions, could also be valuable considerations. Addressing these limitations would help further advance procedural mistake detection and improve its effectiveness in dynamic and diverse environments.

### v3

@@ 5.1.2 Prompt Analysis
The results in Table 4 highlight the positive impact of combining Automatic-Chain-of-Though (ACoT) reasoning with Few-Shot (FS) examples on the performance of LLAMA 3.1 8B across different modalities in the Assembly101-O dataset. In all prompting scenarios, the system prompt is written as:
{systemframe}
System: I am going to provide an input sequence that represents a sequence of actions. Your task is to predict the next action of the last sequence based on the patterns observed in the provided input. Limit yourself to only answer with the predicted sequence, and follow the same format given as input.
@@
In Zero-Shot (ZS) settings, the instruction prompt is presented as:
{systemframe}
System: Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
@@
In the textual modality, where actions are represented by their names (e.g., attach-wheel), adding the system prompt increases $\mathrm{\text{Precision}}_{b}$ from 54.9% to 56.7% (+1.8%), and Recall also rises from 86.5% to 89.6% (+3.1%) , yielding an $\mathrm{\text{F1 score}}_{b}$ gain from 67.1% to 69.4% (+2.3%).
In the numerical modality, where actions are encoded by their index, including the system prompt boosts $\mathrm{\text{Precision}}_{b}$ from 52.4% to 54.5% (+2.1%) and Recall shifts from 94.6% down to 90.3% (–4.3%), with a modest $\mathrm{\text{F1 score}}_{b}$ decrease from 65.7% to 67.9% (+2.2%). The improvements obtained while using the textual representation can be attributed to the system prompt and few-shot examples, which provide explicit context and structural guidance, enabling the model to better understand the task and produce more accurate, relevant predictions.
@@
When comparing ZS, FS, and ACoT prompting methods, ACoT outperforms the rest of the prompting techniques (Table 5), particularly when using the pre-trained model (Base), across both textual and numerical tasks. In the ACoT approach, automatic ACoT reasoning is employed, where the LLM is queried twice: first, to generate an internal reasoning step for the answer, and then to use this reasoning, combined with a few examples, as in FS, to derive the final answer. The first query is:
@@
ACoT, in contrast, excels in these scenarios because it prompts the model to break down the task into a series of logical intermediate steps before concluding (Fig. 6). This approach mirrors the “self-explanation” cognitive strategy, where breaking down and explaining each step improves problem-solving skills. This explains why ACoT significantly outperforms ZS and FS, especially in tasks that demand nuanced understanding and intermediate reasoning.
Figure 6: Visualization of aCoT. First, the aCoT prompt is provided (1) to the LLM to enable (2) the generation of the CoT. The obtained reasoning is then fed to the model again together with the In-Context prompt (3, 4) to retrieve (5) the educated answer from the model.
@@ 5.1.4 Speed Analysis
Table 6 presents a speed test comparison of different prompting methods—ZS, FS, ACoT—using LLAMA 3.1 on the Assembly101-O dataset. The results indicate that ZS is the fastest method, with an average processing speed of 0.208 seconds per sample. FS is slightly slower at 0.216 seconds per sample, representing a marginal 3.8% rise in processing time compared to ZS. ACoT, while offering superior accuracy and reasoning ability, is the slowest, taking 0.315 seconds per sample, which is a 51.4% increase in speed compared to ZS.
@@
These differences in speed can be attributed to the inherent complexity of each method. ZS and FS prompt the model to generate answers directly, with FS introducing a minimal overhead due to the additional context provided by the few examples. On the other hand, ACoT requires the model to engage in more complex, step-by-step reasoning, which naturally demands more computational resources and time. While ACoT’s performance benefits are clear, these come at the cost of slower processing, which could be a consideration in time-sensitive applications. Thus, the choice of prompting method should balance the trade-off between speed and output quality based on the task’s specific requirements.
For completeness, we also report the speed analysis of the language models used in Table 2. The aggregation window is set to $W=200$ frames at 30 fps, corresponding to an effective delay of approximately $200/30\approx 6.67~\text{s}$. For all LLMs (See Table 7), inference latency is negligible by comparison. Thus, when decisions are made after temporal aggregation, the dominant latency stems from the windowing process required to stabilize per-frame recognition, rather than from the LLM inference itself.
Table 7: Speed test on baseline LLM methods.
Model Anticipator speed (samples/s) Total speed (steps/s)
PREGO [Flaborea_2024_CVPR] – GPT-3.5 1.436 8.096
PREGO [Flaborea_2024_CVPR] – DeepSeek-R1 0.612 7.272
PREGO [Flaborea_2024_CVPR] – LLaMA 3.1 0.216 6.876
TI-PREGO – LLaMA 3.1 0.315 6.975
@@ 5.2 Aggregation Strategy
In this section, we quantitatively evaluate the impact of each proposed aggregation strategy (cf. Sec.3.2). As can be seen in Fig.7, all the three aggregation schemes result in a cleaner input for the anticipator module.
@@
We assess the impact of the proposed aggregation strategies using the Levenshtein Similarity between the predicted and ground truth action sequence, which derives from the Levenshtein distance. With this choice, the similarity between two action sequences (i.e. procedures) A and B is defined as:
$\text{Levenshtein Similarity}(A,B)=1-\frac{d_{Lev}(A,B)}{\max(len(A),len(B))}$ (6)
@@ 6 Conclusion
Real-time processing constraints can present challenges, particularly when incorporating Large Language Models (LLMs), which may introduce latency during inference. To clarify the timing budget, our aggregation window of $W=200$ frames at 30 fps induces an effective delay of approximately 6.7s, while the LLM anticipator adds on average $\sim 0.3$,s per query. Hence, the dominant latency arises from the aggregation needed to stabilize per-frame recognition rather than from the LLM inference itself. Our focus is maintaining accuracy in an online setting, rather than optimizing for real-time performance. Nonetheless, as recognition quality improves, smaller windows may suffice, reducing aggregation-induced delay and making real-time optimization a promising direction for future work.
While symbolic reasoning simplifies certain aspects of action prediction, it may not fully capture the complex semantic relationships between actions. In fact, LLMs, when used as in-context learners, can better capture such dependencies through contextual understanding and broad world knowledge. However, their reasoning process is not explicitly based on logic rules and relies on patterns learned from data, which may limit their ability to model fine-grained or domain-specific action dependencies. Future research could explore hybrid approaches that integrate symbolic reasoning with richer representations.
Our robustness treatment is pragmatic: aggregation reduces short-horizon fluctuations so that downstream anticipation operates on a steadier signal. Ambiguous cases, like brief transitions or rapidly alternating labels, are mitigated but not eliminated by this design. We use ACoT’s intermediate explanations as qualitative aids to inspect typical successes and failures. These examples improve transparency but do not constitute a formal faithfulness guarantee. A systematic taxonomy of ambiguity and formal uncertainty quantification are valuable directions for future work and lie beyond the scope of this manuscript.
@@ Acknowledgements
AI Assistance Declaration
During the preparation of this work, the authors used Large Language Models to assist in formulating the text. After using these tools, the authors carefully reviewed and edited the content as necessary and take full responsibility for the final version of the published article.